Operations Report – April 7th

Crew 279 Operations Report 07-04-2023

SOL: 5

Name of person filing report: Ioana Dimitrova

Non-nominal systems: robotic observatory, Starlink network outage, Opportunity

Notes on non-nominal systems:
Robotic observatory: awaiting for a solution from the manufacturer.
Starlink outage: see Internet section.
Opportunity Rover : see rover section

Spirit rover used: no
Hours: 214,2
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Opportunity rover used: yes
Hours: 120
Beginning charge: 100%
Ending charge: 77%
Currently charging: yes

Curiosity rover used: no
Hours: 227.1
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Perseverance rover used: yes
Hours: 260.3
Beginning charge: 100%
Ending charge: 77%
Currently charging: yes

General notes on rovers: one of the battery wires on the driver side of Opportunity is corroded out and needs to be replaced.

Summary of Hab operations: Music influence on stress experiment was conducted. Screws on suit 8 were tightened as they were coming a bit loose. Data on all available food was collected for the microbiota experiment.
Water (static tank): 389,9
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet: used from 14:00 to 21:00 for experiments and reports. Starlink outage duration 22m 47 sec, followed by significant internet speed drop.

Summary of suits and radios: Radios charging nominal. Suits charging nominal. Screws were tightened on suit 8 as they were a bit loose. Suit 8 (the one with the new battery) was out on EVA today, everything nominal.

WATER USE: 10.56
Heater: Off
Supplemental light: Off
Harvest: 40g mint, 10g lemon balm, 15g thyme, 13g oregano, 19g sage

Summary of ScienceDome operations : bacteria cultures were packaged for balloon experiment. Filming and editing for mid-mission video.

Dual split: Heat or AC, Off

Summary of RAM operations: 3D design and test printing of new and better knobs for the oven following Operations Report of SOL 1 (stem too long to turn in an optimal way). If they work better than old knobs we will ask Mission support if we can replace the old ones. Old knobs still work nominal but could be optimized.

Summary of any observatory issues: see above for robotic observatory. Sun observatory opened but no photos because of cloudy skies. See observatory report for more details.

Summary of health and safety issues: none

Questions, concerns and requests to Mission Support:
We were asked to use power to max in the morning and return to normal in the afternoon, following yesterday’s power outage.

Copyright © The Mars Society. All rights reserved. | Main Site